In San Francisco, 30% of workers take public transportation daily.
a. In a sample of ten workers. what is the probability that exactly three workers take public transportation daily?
b. In a sample of 10 workers, what is the probability that at least three workers take public transportation daily?

Answer by edjones(8007)


a)
x=.3, y=.7
10C3=120 Binomial Theorem
120*x^3*y^7=.267
.
b)
10C2=45
45*x^2*y^8=.234
.
10C1=10
10*x*y^9=.121
.
10C0=1
y10=.7^10=.028
.
.234+.121+.028=.383 Probability of less than 3 workers taking public trans.
1-.383=.617 the probability that at least three workers take public transportation daily.
.
Ed